OpenStudy (anonymous):

Another one @ganeshie8

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}\frac{ 1 }{ \sum_{n=1}^{\infty} k^3}\]

ganeshie8 (ganeshie8):

the index in denominator is \(k\) right ?

ganeshie8 (ganeshie8):

\[\sum_{n=1}^{\infty}\frac{ 1 }{ \sum_{\color{red}{k}=1}^{\infty} k^3}\]

OpenStudy (anonymous):

YUP SRY

OpenStudy (anonymous):